Solve for x. |2x-4|+1=x
Answer by stanbon(75887)

Solve for x. |2x-4|+1=x
------
|2x-4| = x-1
---
2x-4 = x-1 or 2x-4 = -(x-1)
x = 3 or x = 5
======================
Cheers,
Stan H.